A mobile game is such that the character will always walk 3 steps North, 4 steps East, 11 steps

South and 10 steps West in the same order, over and over again. How far is the character from

its original position after doing the procedure 20 times?


1
Expert's answer
2021-01-22T14:21:26-0500

Position after 1st procedure:

11-3=8 steps South

10-4=6 steps West.

Position after 20th procedure:

8*20=160 steps South

6*20=120 steps West.